A solution of ammonia and water contains 4.20×1025 water molecules and 7.30×1024 ammonia molecules. How many total hydrogen atoms are in this solution?
1
Expert's answer
2015-07-09T06:00:14-0400
1 H2O molecule contain 2 H atoms:
H atoms in 4.20*10^25 water molecules = 2* 4.20 * 10^25 = 8.4 * 10^25 H atoms

1 NH3 molecule contains 3 H atoms:
H atoms in 6.00 * 10^24 NH3 atoms = 3 * 6.00 * 10^24 = 1.8 * 10^25 H atoms

Total H atoms = 8.4 * 10^25 + 1.8 * 10^25 = 1.02 * 10^16 H atoms.